# 2Universität Heidelberg, Zentrum für Astronomie (ZAH), Institut für Theoretische Astrophysik, Albert Ueberle Str. 2, 69120, Heidelberg, Germany 3Universität Heidelberg, Interdiszipliäres Zentrum für Wissenschaftliches Rechnen, D-69120 Heidelberg, Germany 4Dipartimento di Fisica e Astronomia Galileo Galilei, Università di Padova, Vicolo dell’Osservatorio 3, I–35122 Padova, Italy

### v1

@@ 2.4 Population-synthesis with sevn
We explore the effect of optically thick winds on the mass spectrum of BHs, BBHs and BBH mergers using the population-synthesis code sevn (Spera & Mapelli 2017; Spera et al. 2019; Mapelli et al. 2020; Iorio et al. 2023) [^2] We used version 2.10.1 of sevn, updated to commit c9dc8e457. sevn is publicly available here.. sevn interpolates pre-computed stellar tracks to evolve single and binary stars. Also, it incorporates semi-analytic formulas for processes like binary mass transfer, tides, supernova explosion, and gravitational-wave decay. We refer to Iorio et al. (2023) for a complete description of the code.
@@
We generate new evolutionary tables for stars with mass $M_{\star}\geq 50\,{}\mathrm{M}_{\odot}$ from the aforementioned VMS models with mesa. We then evolve single and binary stars with primary mass in this mass range. For binary stars, we use the stellar tables from parsec (Bressan et al. 2012; Chen et al. 2015; Costa et al. 2019a, b; Nguyen et al. 2022; Costa et al. 2025) to account for the evolution of secondary stars with mass $M_{\star}<50\,{\rm M}_{\odot}$.
@@
When two stars collide, we use the standard formalism of sevn as described by Iorio et al. (2023). Specifically, we assume that the collision product retains the total mass of the two stars. The CO and He core masses of the collision product are the sum of the CO and He core masses of the two colliding stars. Finally, the collision product inherits the phase and percentage of life of the most evolved progenitor star. Hence, our models likely overestimate the total mass of the collision product and underestimate mixing. For a more self-consistent treatment of massive star collisions, see, e.g., Costa et al. (2022b) and Ballone et al. (2023).
@@
We randomly draw the zero-age main sequence (ZAMS) primary star masses from a Kroupa (2001) initial mass function, with $M_{\mathrm{ZAMS}}\in\ [50,500]$. For binaries, we draw the secondary masses from the observation-based mass ratio distribution $\mathcal{F}(q)\propto q^{-0.1},\;\mathrm{with}\;q\;\in[0.1,\,{}1]$ from Sana et al. (2012). We generate the initial orbital periods ($P$) and eccentricities ($e$) from the distributions by Sana et al. (2012): $\mathcal{F}(\mathcal{P})\propto\mathcal{P}^{-0.55},\;\mathrm{with}\;\mathcal{P}=\log_{10}(P/{\rm days})\in[0.15,\,{}5.5]$ and $\mathcal{F}(e)\propto e^{-0.45},\;\mathrm{with}\;e\in[10^{-5},\,{}e_{\rm max}(P)]$, respectively. For a given orbital period, we set the upper limit of the eccentricity distribution $e_{\rm max}(P)$ as $e_{\rm max}(P)=1-\left[P/(2\,{\rm days})\right]^{-2/3}$ (Moe & Di Stefano 2017).
We investigate eight different metallicities: $Z=0.0001$, 0.001, 0.002, 0.004, 0.006, 0.008, 0.01, 0.02. For each metallicity, we generate $2\times 10^{6}$ single and $1\times 10^{7}$ binary systems. Our initial setup corresponds to the same as the fiducial model by Iorio et al. (2023). To highlight the impact of stellar winds on the resulting BH mass distribution, we also perform a run with tables from parsec tracks only. In this case, the stellar wind prescription is purely based on the wind model introduced by Chen et al. (2015, hereafter C15), which, in turn is derived from Vink et al. (2000, 2001) for the winds of O-type stars, with additional corrections for electron scattering, the mass-loss dependence on the Eddington ratio (Gräfener & Hamann 2008; Vink et al. 2011). We refer to Simonato et al. (2025) for a thorough comparison between the two wind prescriptions.
@@ 3.1 Stellar cores
Figure 1 shows the final He and CO core masses for the S23 and C15 wind models. At low metallicity ($Z\leq{}0.0001$), the two wind models predict the same monotonic increase of the He and CO core mass with $M_{\mathrm{ZAMS}}$, because VMSs do not enter the optically thick wind regime. At these metallicities, the C15 models generally yield slightly smaller cores, due to the different envelope undershooting and convection models (Simonato et al. 2025).
At higher metallicity, the two considered models C15 and S23 behave in a dramatically different way. According to S23, VMSs can enter the optically thick wind regime at $Z\geq{}0.001$: such strong winds reduce the final total mass and core mass of VMSs, resulting in a flat or even decreasing trend of $M_{\rm c}$ with $M_{\rm ZAMS}$. Hence, the S23 model predicts smaller He and CO cores compared to C15, with a deep impact on the occurrence of (pulsational) pair instability.
The effect of optically thick winds in the S23 model is twofold. On the one hand, the S23 model results in a lower metallicity threshold below which VMSs can enter the (pulsational) pair instability regime: the threshold is $Z_{\rm th}=0.004$ and $0.01$ for S23 and C15, respectively. On the other hand, the formation of IMBHs via direct collapse is suppressed in the S23 model for metallicity $Z\geq{}0.001$, whereas the C15 model allows the formation of IMBHs already at $Z\leq{}0.008$ because of the lower wind mass-loss.
For example, at $Z=0.001$ optically thick winds become effective for $M_{\mathrm{ZAMS}}>200\,{\rm M}_{\odot}$, when the stellar core is already well within the pair-instability regime. Mass loss prevents the He core mass from increasing above $135\,{\rm M}_{\odot}$. As a consequence, only VMSs with $M_{\mathrm{ZAMS}}\gtrsim 450\,{\rm M}_{\odot}$ undergo direct collapse to IMBH in the S23 model, whereas in C15 IMBHs can form via direct collapse already at $M_{\mathrm{ZAMS}}\sim 300\,{\rm M}_{\odot}$.
@@
For $Z\gtrsim 0.006$, VMSs in S23 models undergo optically thick stellar winds throughout the mass range considered. This dramatically reduces final core masses, which display a weak dependence on $M_{\mathrm{ZAMS}}$. For example, at $Z\sim 0.006$ the core mass for VMSs with $M_{\mathrm{ZAMS}}>400\,{\rm M}_{\odot}$ decreases from $\sim 200\,{\rm M}_{\odot}$ in the C15 model to $\sim 20\,{\rm M}_{\odot}$ in S23.
@@ 3.2 BH masses
Figure 2 shows the final BH mass as a function of $M_{\mathrm{ZAMS}}$ for models C15 and S23. The key difference between the two models is the metallicity threshold to form IMBHs; this is $Z_{\rm IMBH}=0.001$ and 0.012 for models S23 and C15, respectively. In model S23, IMBH formation is suppressed at metallicity Z>0.001Z>{0.001} by the onset of optically thick winds. These strong winds erode the mass of the He core and prevent it from reaching the threshold $M_{\rm He}\sim{135}$ M, above which the star avoids to explode as a pair-instability supernova. Thus, even the most massive stars do not produce IMBHs in the S23 model unless they have metallicity $Z\leq{}0.001$. In contrast, the lower mass loss in model C15 allows to grow larger cores and enables the formation of IMBHs up to $Z\sim{0.012}$.
@@
In absence of optically thick winds, the BH mass increases with $M_{\mathrm{ZAMS}}$, up to the onset of (pulsational) pair instability, at $M_{\mathrm{ZAMS}}\gtrsim 75\,{\rm M}_{\odot}$. For larger initial masses, the onset of direct collapse depends on metallicity. In the C15 model, VMSs undergo (pulsational) pair instability up to $Z=0.01$. At lower metallicities, all the models with $M_{\mathrm{ZAMS}}>400\,{\rm M}_{\odot}$ produce IMBHs, with masses ranging from $135\,{\rm M}_{\odot}$ to $450\,{\rm M}_{\odot}$. At $Z=0.02$, instead, wind mass-loss is strong enough to prevent the onset of pair instability, leading to a maximum BH mass $M_{\mathrm{BH}}\sim 30\,{\rm M}_{\odot}$.
@@
In the S23 model, the maximum metallicity at which IMBHs form is an order of magnitude lower ($Z=0.001$) compared to C15. Below this threshold, the S23 model produces IMBHs with masses in the same range as C15. At higher metallicities, the most massive BHs originate from stars with $M_{\mathrm{ZAMS}}\sim{50}{\rm M}_{\odot}$, while more massive stars deliver $M_{\mathrm{BH}}\lesssim 15{\rm M}_{\odot}$. This happens because VMSs develop optically thick winds at an early stage, and spend all their lifetime with $\dot{M}>10^{-5}$ M yr-1. This results in a severe erosion of the core (Simonato et al. 2025; Boco et al. 2025). Overall, the S23 models predict a pair-instability BH mass gap from $65\,{\rm M}_{\odot}$ to $135\,{\rm M}_{\odot}$.
@@
Figure 3 shows the final BH mass as a function of the ZAMS mass and the metallicity for the S23 model. The formation of IMBHs takes place only for VMSs with $M_{\mathrm{ZAMS}}\gtrsim 250\,{\rm M}_{\odot}$ at $Z\leq 0.001$. Also, the formation of BHs with $M_{\mathrm{BH}}>20\,{}{\rm M}_{\odot}$ is suppressed for $Z>0.01$ throughout the mass range.
@@
Figure 4: BH mass distribution from single stars with $M_{\mathrm{ZAMS}}>50\,{}{\rm M}_{\odot}$, for the models adopting the S23 (upper panel) and C15 (lower panel) wind prescriptions and for different metallicities.
@@
Figure 5: BH mass distribution from binaries with $M_{\mathrm{ZAMS,1}}>50{\rm M}_{\odot}$, for models adopting the S23 (blue, solid line, solid fill) and C15 (orange, dashed line, hatched area) wind prescriptions. The shaded (hatched) areas indicate the BHs in BBHs, while the empty histograms display the single BHs from disrupted binaries, including BHs that form from stellar collisions.
@@ 3.3.1 BHs from single VMSs
Figure 4 shows the BH mass function from single VMSs assuming a Kroupa (Kroupa 2001) initial mass function. For $Z>0.001$ optically thick winds strongly suppress the production of IMBHs from single stars. At $Z=0.001$, the S23 model still allows IMBH formation, but the efficient mass loss limits the maximum BH mass to $150\,{\rm M}_{\odot}$, compared to $350\,{\rm M}_{\odot}$ in the C15 model. For higher metallicities, stellar winds progressively reduce the maximum BH mass in the S23 model, down to $20\,{\rm M}_{\odot}$ at $Z=0.02$.
In the C15 model, the formation of IMBHs takes place up to $Z=0.01$, where the maximum BH mass is $\sim 180\,{\rm M}_{\odot}$ (see also Costa et al. 2025). At lower metallicities, the maximum IMBH mass reaches $\sim{400}$ M, for ZAMS masses up to $_{\rm ZAMS}=500 M$. In this model, IMBH formation from single VMSs is only suppressed at $Z=0.02$, where BHs can form with a mass up to $33\,{\rm M}_{\odot}$.
@@ 3.3.2 BBHs from VMS binaries
The C15 model predicts the formation of BHs as massive as $\gtrsim 700\,{\rm M}_{\odot}$ for $Z\leq 0.002$. Also, IMBHs up to $\sim 400\,{\rm M}_{\odot}$ can still form at $Z=0.01$. In the S23 model, instead, the formation of IMBHs is suppressed already at $Z=0.004$, where they represent less than 0.2% of the total BH number, despite the occurrence of star-star colllisions. In contrast, the percentage of IMBHs is $21\%$ in the C15 model at $Z=0.004$.
Primary BHs in BBHs also reach higher masses than BHs from single stellar evolution, due to mass accretion from the stellar companion. In the S23 model, IMBH formation in BBHs is possible for $Z\leq 0.002$, while for $Z=0.004$, the maximum primary BH mass in BBHs decreases to $45\,{\rm M}_{\odot}$. At $Z=0.02$, the BH mass function in the S23 model extends from $5\,{\rm M}_{\odot}$ to $20\,{\rm M}_{\odot}$.
BHs in BBHs display a pair-instability mass gap between $65\,{\rm M}_{\odot}$ and $135\,{\rm M}_{\odot}$ for both the C15 and the S23 models. In contrast, star-star collisions can produce BHs in this mass range, and fill the pair-instability mass gap with single BHs.

### v2

@@ 2.4 Population-synthesis with sevn
We explore the effect of optically thick winds on the mass spectrum of BHs, BBHs and BBH mergers using the population-synthesis code sevn (Spera2017; Spera2019; Mapelli2020; Iorio2023) [^3] We used version 2.10.1 of sevn, updated to commit c9dc8e457. sevn is publicly available here.. sevn interpolates pre-computed stellar tracks to evolve single and binary stars. Also, it incorporates semi-analytic formulas for processes like binary mass transfer, tides, supernova explosion, and gravitational-wave decay. We refer to Iorio2023 for a complete description of the code.
@@
We generate new evolutionary tables for stars with mass $M_{\star}\geq 50\,{}\mathrm{M}_{\odot}$ from the aforementioned VMS models with mesa. We then evolve single and binary stars with primary mass in this mass range. For binary stars, we use the stellar tables from parsec (Bressan2012; Chen2015; Costa2019b; Costa2019a; Nguyen2022; costa2025) to account for the evolution of secondary stars with mass $M_{\star}<50\,{\rm M}_{\odot}$.
@@
When two stars collide, we use the standard formalism of sevn as described by Iorio2023. Specifically, we assume that the collision product retains the total mass of the two stars. The CO and He core masses of the collision product are the sum of the CO and He core masses of the two colliding stars. Finally, the collision product inherits the phase and percentage of life of the most evolved progenitor star. Hence, our models likely overestimate the total mass of the collision product and underestimate mixing. For a more self-consistent treatment of massive star collisions, see, e.g., costa2022a and Ballone2023.
@@
We randomly draw the zero-age main sequence (ZAMS) primary star masses from a Kroupa2001 initial mass function, with $M_{\mathrm{ZAMS}}\in\ [50,500]$. For binaries, we draw the secondary masses from the observation-based mass ratio distribution $\mathcal{F}(q)\propto q^{-0.1},\;\mathrm{with}\;q\;\in[0.1,\,{}1]$ from sana2012. We generate the initial orbital periods ($P$) and eccentricities ($e$) from the distributions by sana2012: $\mathcal{F}(\mathcal{P})\propto\mathcal{P}^{-0.55},\;\mathrm{with}\;\mathcal{P}=\log_{10}(P/{\rm days})\in[0.15,\,{}5.5]$ and $\mathcal{F}(e)\propto e^{-0.45},\;\mathrm{with}\;e\in[10^{-5},\,{}e_{\rm max}(P)]$, respectively. For a given orbital period, we set the upper limit of the eccentricity distribution $e_{\rm max}(P)$ as $e_{\rm max}(P)=1-\left[P/(2\,{\rm days})\right]^{-2/3}$ (moe2017).
We investigate eight different metallicities: $Z=0.0001$, 0.001, 0.002, 0.004, 0.006, 0.008, 0.01, 0.02. For each metallicity, we generate $2\times 10^{6}$ single and $1\times 10^{7}$ binary systems. Our initial setup corresponds to the same as the fiducial model by Iorio2023.
To highlight the impact of stellar winds on the resulting BH mass distribution, we also perform a run with tables from the non-rotating parsec tracks presented by costa2025444The parsec stellar tracks are publicly available at http://stev.oapd.inaf.it/PARSEC/ (costa2025).. In this case, the adopted wind model is the same as described by Chen2015. Specifically, the winds of massive O-type stars are modeled adopting the fitting formulas by Vink2000 and Vink2001, with the correction by Graefener2008 to account for the effects of electron scattering. The mass-loss rate of an O-type star scales with the metallicity as $\dot{M}\propto Z^{\beta}$, where $\beta$ depends on the Eddington ratio $\Gamma_{\rm Edd}$: $\beta=0.85$ for $\Gamma_{\rm Edd}<2/3$, $\beta=2.45-2.4\times\Gamma_{\rm Edd}$ for $2/3\leq\Gamma_{\rm Edd}<1$, and $\beta=0.05$ for $\Gamma_{\rm Edd}\geq 1$ (Chen2015). In this way, this formalism incorporates the dependence of mass loss on the Eddington ratio (Graefener2008; Vink2011). As for Wolf–Rayet (WR) stars, parsec adopts the prescriptions by Sander2019, which match the properties of Galactic WR stars of type C (WC) and type O (WO). We refer to costa2025 for other details on these stellar tracks.
@@ 3.1 Stellar cores
Figure 1 shows the final He and CO core masses for the Sabhahit2023 and Chen2015 wind models. At low metallicity ($Z\leq{}0.0001$), the two wind models predict the same monotonic increase of the He and CO core mass with $M_{\mathrm{ZAMS}}$, because VMSs do not enter the optically thick wind regime. At these metallicities, the Chen2015 models generally yield slightly smaller cores, due to the different envelope undershooting and convection models (simonato2025).
At higher metallicity, the two considered models Chen2015 and Sabhahit2023 behave in a dramatically different way. According to Sabhahit2023, VMSs can enter the optically thick wind regime at $Z\geq{}0.001$: such strong winds reduce the final total mass and core mass of VMSs, resulting in a flat or even decreasing trend of $M_{\rm c}$ with $M_{\rm ZAMS}$. Hence, the Sabhahit2023 model predicts smaller He and CO cores compared to Chen2015, with a deep impact on the occurrence of (pulsational) pair instability.
The effect of optically thick winds in the Sabhahit2023 model is twofold. On the one hand, the Sabhahit2023 model results in a lower metallicity threshold below which VMSs can enter the (pulsational) pair instability regime: the threshold is $Z_{\rm th}=0.004$ and $0.01$ for Sabhahit2023 and Chen2015, respectively. On the other hand, the formation of IMBHs via direct collapse is suppressed in the Sabhahit2023 model for metallicity $Z\geq{}0.001$, whereas the Chen2015 model allows the formation of IMBHs already at $Z\leq{}0.008$ because of the lower wind mass-loss.
For example, at $Z=0.001$ optically thick winds become effective for $M_{\mathrm{ZAMS}}>200\,{\rm M}_{\odot}$, when the stellar core is already well within the pair-instability regime. Mass loss prevents the He core mass from increasing above $135\,{\rm M}_{\odot}$. As a consequence, only VMSs with $M_{\mathrm{ZAMS}}\gtrsim 450\,{\rm M}_{\odot}$ undergo direct collapse to IMBH in the Sabhahit2023 model, whereas in Chen2015 IMBHs can form via direct collapse already at $M_{\mathrm{ZAMS}}\sim 300\,{\rm M}_{\odot}$.
@@
For $Z\gtrsim 0.006$, VMSs in Sabhahit2023 models undergo optically thick stellar winds throughout the mass range considered. This dramatically reduces final core masses, which display a weak dependence on $M_{\mathrm{ZAMS}}$. For example, at $Z\sim 0.006$ the core mass for VMSs with $M_{\mathrm{ZAMS}}>400\,{\rm M}_{\odot}$ decreases from $\sim 200\,{\rm M}_{\odot}$ in the Chen2015 model to $\sim 20\,{\rm M}_{\odot}$ in Sabhahit2023.
@@ 3.2 BH masses
Figure 2 shows the final BH mass as a function of $M_{\mathrm{ZAMS}}$ for models Chen2015 and Sabhahit2023. The key difference between the two models is the metallicity threshold to form IMBHs; this is $Z_{\rm IMBH}=0.001$ and 0.012 for models Sabhahit2023 and Chen2015, respectively. In model Sabhahit2023, IMBH formation is suppressed at metallicity Z>0.001Z>{0.001} by the onset of optically thick winds. These strong winds erode the mass of the He core and prevent it from reaching the threshold $M_{\rm He}\sim{135}$ M, above which the star avoids to explode as a pair-instability supernova. Thus, even the most massive stars do not produce IMBHs in the Sabhahit2023 model unless they have metallicity $Z\leq{}0.001$. In contrast, the lower mass loss in model Chen2015 allows to grow larger cores and enables the formation of IMBHs up to $Z\sim{0.012}$.
@@
In absence of optically thick winds, the BH mass increases with $M_{\mathrm{ZAMS}}$, up to the onset of (pulsational) pair instability, at $M_{\mathrm{ZAMS}}\gtrsim 75\,{\rm M}_{\odot}$. For larger initial masses, the onset of direct collapse depends on metallicity. In the Chen2015 model, VMSs undergo (pulsational) pair instability up to $Z=0.01$. At lower metallicities, all the models with $M_{\mathrm{ZAMS}}>400\,{\rm M}_{\odot}$ produce IMBHs, with masses ranging from $135\,{\rm M}_{\odot}$ to $450\,{\rm M}_{\odot}$. At $Z=0.02$, instead, wind mass-loss is strong enough to prevent the onset of pair instability, leading to a maximum BH mass $M_{\mathrm{BH}}\sim 30\,{\rm M}_{\odot}$.
@@
In the Sabhahit2023 model, the maximum metallicity at which IMBHs form is an order of magnitude lower ($Z=0.001$) compared to Chen2015. Below this threshold, the Sabhahit2023 model produces IMBHs with masses in the same range as Chen2015. At higher metallicities, the most massive BHs originate from stars with $M_{\mathrm{ZAMS}}\sim{50}{\rm M}_{\odot}$, while more massive stars deliver $M_{\mathrm{BH}}\lesssim 15{\rm M}_{\odot}$. This happens because VMSs develop optically thick winds at an early stage, and spend all their lifetime with $\dot{M}>10^{-5}$ M yr-1. This results in a severe erosion of the core (simonato2025; boco2025). Overall, the Sabhahit2023 models predict a pair-instability BH mass gap from $65\,{\rm M}_{\odot}$ to $135\,{\rm M}_{\odot}$.
@@
Figure 3 shows the final BH mass as a function of the ZAMS mass and the metallicity for the Sabhahit2023 model. The formation of IMBHs takes place only for VMSs with $M_{\mathrm{ZAMS}}\gtrsim 250\,{\rm M}_{\odot}$ at $Z\leq 0.001$. Also, the formation of BHs with $M_{\mathrm{BH}}>20\,{}{\rm M}_{\odot}$ is suppressed for $Z>0.01$ throughout the mass range.
@@
Figure 4: BH mass distribution from single stars with $M_{\mathrm{ZAMS}}>50\,{}{\rm M}_{\odot}$, for the models adopting the Sabhahit2023 (upper panel) and Chen2015 (lower panel) wind prescriptions and for different metallicities.
@@
Figure 5: BH mass distribution from binaries with $M_{\mathrm{ZAMS,1}}>50{\rm M}_{\odot}$, for models adopting the Sabhahit2023 (blue, solid line, solid fill) and Chen2015 (orange, dashed line, hatched area) wind prescriptions. The shaded (hatched) areas indicate the BHs in BBHs, while the empty histograms display the single BHs from disrupted binaries, including BHs that form from stellar collisions.
@@ 3.3.1 BHs from single VMSs
Figure 4 shows the BH mass function from single VMSs assuming a Kroupa (Kroupa2001) initial mass function. For $Z>0.001$ optically thick winds strongly suppress the production of IMBHs from single stars. At $Z=0.001$, the Sabhahit2023 model still allows IMBH formation, but the efficient mass loss limits the maximum BH mass to $150\,{\rm M}_{\odot}$, compared to $350\,{\rm M}_{\odot}$ in the Chen2015 model. For higher metallicities, stellar winds progressively reduce the maximum BH mass in the Sabhahit2023 model, down to $20\,{\rm M}_{\odot}$ at $Z=0.02$.
In the Chen2015 model, the formation of IMBHs takes place up to $Z=0.01$, where the maximum BH mass is $\sim 180\,{\rm M}_{\odot}$ (see also costa2025). At lower metallicities, the maximum IMBH mass reaches $\sim{400}$ M, for ZAMS masses up to $_{\rm ZAMS}=500 M$. In this model, IMBH formation from single VMSs is only suppressed at $Z=0.02$, where BHs can form with a mass up to $33\,{\rm M}_{\odot}$.
@@ 3.3.2 BBHs from VMS binaries
The Chen2015 model predicts the formation of BHs as massive as $\gtrsim 700\,{\rm M}_{\odot}$ for $Z\leq 0.002$. Also, IMBHs up to $\sim 400\,{\rm M}_{\odot}$ can still form at $Z=0.01$. In the Sabhahit2023 model, instead, the formation of IMBHs is suppressed already at $Z=0.004$, where they represent less than 0.2% of the total BH number, despite the occurrence of star-star colllisions. In contrast, the percentage of IMBHs is $21\%$ in the Chen2015 model at $Z=0.004$. We find that, in the metallicity range Z=0.0040.008Z=0.004-0.008, our distributions are consistent with those from shepherd2025 when they consider the same wind prescription.
Primary BHs in BBHs also reach higher masses than BHs from single stellar evolution, due to mass accretion from the stellar companion. In the Sabhahit2023 model, IMBH formation in BBHs is possible for $Z\leq 0.002$, while for $Z=0.004$, the maximum primary BH mass in BBHs decreases to $45\,{\rm M}_{\odot}$. At $Z=0.02$, the BH mass function in the Sabhahit2023 model extends from $5\,{\rm M}_{\odot}$ to $20\,{\rm M}_{\odot}$.
BHs in BBHs display a pair-instability mass gap between $65\,{\rm M}_{\odot}$ and $135\,{\rm M}_{\odot}$ for both the Chen2015 and the Sabhahit2023 models. In contrast, star-star collisions can produce BHs in this mass range, and fill the pair-instability mass gap with single BHs.